OpenStudy (anonymous):

-2x/7-3/7 i know how to do it but im just to lazy to do it. and this isnt really for a grade either so.

OpenStudy (kc_kennylau):

\[\frac{-2x-3}7\]

OpenStudy (kc_kennylau):

Is this an equation

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

-.- you act like im really stupid

OpenStudy (kc_kennylau):

is. this. an. equation.

OpenStudy (kc_kennylau):

i know u ain't stupid

OpenStudy (anonymous):

there is no equal sign

OpenStudy (anonymous):

wowza finals makes everyone lazy

OpenStudy (kc_kennylau):

@kelliegirl33 i'm asking if this is an equation in case he missed out the equal sign somewhere

OpenStudy (anonymous):

yup and *she*

OpenStudy (anonymous):

equations have equal signs i thought

OpenStudy (anonymous):

\[\frac{ -2x }{ 7}-\frac{ 3 }{ 7 }= ?\]

OpenStudy (kc_kennylau):

@kelliegirl33 because i don't find any point in \(\dfrac{-2x-3}7\)

OpenStudy (anonymous):

this always turns into a big deal

OpenStudy (anonymous):

its not were all just confused as too what you are asking for

OpenStudy (anonymous):

hows it a big deal? lol we're just trying to figure out what exactly you are asking.

OpenStudy (anonymous):

just never mind xD sorry if im confusing

OpenStudy (anonymous):

-2x/7 - 3/7 = ? -2x/7 = 3/7 (multiply by 7) -2x = 3 x = -3/2 check.. -2x/7 - 3/7 = ? -2x/7 = 3/7 (-2(-3/2))/7 = 3/7 (6/2)/7 = 3/7 3/7 = 3/7 (correct) x = -3/2 you will need to get a second opinion on this....I am not even sure if this is right